Could somebody please help me solve 11+4x=21-x
I'm sorry I have already asked this, it's just that nobody answered it.
Thanks!

Answer by mananth(16946)

11+4x=21-x
Add x to both sides
11+4x+x= 21-x+x
11+5x=21
add -11 to both sides
11-11+5x=21-11
5x=10
divide by 5
x=2